# Small Gradient Norm Regret for Online Convex Optimization

### v2

@@ 3.1 Dynamic $G^{\star}$ regret
Section˜2 focuses on the static regret with respect to a single competitor $x\in\mathcal{X}$. In this subsection, we consider dynamic regret, which allows for a sequence of competitors at the cost of a path length term in the final regret bound. Given a comparator sequence $\{\hat{x}_{t}\}_{t=1}^{T}$ with path length $\mathcal{P}_{T}(\{\hat{x}_{t}\}_{t=1}^{T})\coloneqq\sum_{t=2}^{T}\|\hat{x}_{t+1}-\hat{x}_{t}\|$, the dynamic regret with respect to this sequence is defined as
@@ 3.2 $G^{\star}$ regret under bandit feedback
This smoothed function $\tilde{\ell}_{t}$ naturally serves as a surrogate function when performing BGD. Key properties of the smoothed function used in our analysis are summarized in the appendix. The following theorems establish regret bounds for BGD under constant and adaptive learning rates, respectively.
@@
Under A1, A2 and suppose $n\geq 8$, bandit gradient descent (Algorithm˜5 in the appendix) $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta g_{t}]$ with constant learning rate $\eta\in(0,\tfrac{1}{4nL})$ satisfies $\mathbb{E}[\rho_{T}(x)]\leq\textstyle\tfrac{D^{2}}{2\eta}+\tfrac{4n\eta}{1-4n\eta L}G_{T}(x)+\tfrac{4n\eta}{1-4n\eta L}TL^{2}\mu^{2}+\tfrac{\eta T}{2}n^{2}L^{2}\mu^{2}+\tfrac{TL\mu^{2}}{2}$ for any $x\in\mathcal{X}$. In particular, choosing $x=x^{\star}$, $\eta=\min\big\{\tfrac{D}{4\sqrt{nG^{\star}_{T}}},\tfrac{1}{8nL}\big\}$ and $\mu=\tfrac{D}{\sqrt{2nT}L}$ gives $\mathbb{E}[\rho_{T}(x^{\star})]\leq\max\{8nLD^{2},4D\sqrt{nG^{\star}_{T}}\}+\tfrac{LD^{2}}{2}$.
@@ 4.1 Stochastic optimization under interpolation
$\sigma_{f}^{\star}\coloneqq\inf_{x}\mathbb{E}[f(x,\xi)-\inf_{x}f(x,\xi)]\quad\text{and}\quad\sigma_{g}^{\star}\coloneqq\sqrt{\mathbb{E}[\|\nabla f(x^{\star},\xi)\|^{2}]},$
where $x^{\star}\in\operatornamewithlimits{arg\,min}_{x}f(x)$. We also make the following simplifying assumption.
@@ 1. A3:
Under A3, we have $\sigma_{f}^{\star}=\min_{x\in\mathcal{X}}\mathbb{E}[f(x,\xi)-\inf_{x}f(x,\xi)]$, and it is easy to verify that $\sigma_{f}^{\star}$ and $\sigma_{g}^{\star}$ mirror the $L^{\star}$ and $G^{\star}$ regret, respectively. When $\sigma_{f}^{\star}$ or $\sigma_{g}^{\star}$ is close to 0, the regime is known as (near-)interpolation [22]. Interpolation is a central setting in optimization for machine learning and has recently attracted significant attention [30, 21, 34, 2]. Since online algorithms with sublinear regret can be converted into an optimization algorithm via online-to-batch conversion [26], our $G^{\star}$ regret bounds from Section˜2.2 imply convergence guarantees for stochastic gradient methods and stochastic dual averaging (i.e., follow-the-regularized-leader for linearized losses).
@@
  • if OGD from Theorem˜2.1 is applied to $\{\ell_{t}\}_{t=1}^{T}$, then for any $\eta\in(0,\tfrac{1}{L})$, we have
@@
  • if AdaGrad-Norm from Theorem˜2.2 is applied to $\{\ell_{t}\}_{t=1}^{T}$, then $\mathbb{E}[f(\bar{x}^{T})-f(x^{\star})]\leq\tfrac{LD^{2}}{T}+\tfrac{\sqrt{2}D}{\sqrt{T}}\sigma_{g}^{\star}$.
@@
Remark 7.
Corollary˜4.1 improves the existing results [30] by replacing $\sigma_{f}^{\star}$ by $\sigma_{g}^{\star}$. Compared with the recent last-iterate convergence results of stochastic gradient descent [2, 12], our result enables adaptivity to smoothness for the average iterate (Table˜1).
Table 1: Implications of the $G^{\star}$ regret for stochastic smooth convex optimization
@@
[30] Average iterate, $\mathcal{X}=\mathbb{R}^{n}$ SGD No $\mathcal{O}(\tfrac{1}{T}+\tfrac{1}{\sqrt{T}}\sigma_{f}^{\star})$
[2, 12] Last iterate, $\mathcal{X}=\mathbb{R}^{n}$ SGD No $\mathcal{O}\big(\tfrac{\log T}{T}+\sqrt{\tfrac{\log T}{T}}\sigma_{g}^{\star}\big)$
[33] Average iterate, $\operatorname{diam}(\mathcal{X})\leq D$ AdaGrad Yes $\mathcal{O}(\tfrac{1}{T}+\tfrac{1}{\sqrt{T}}\sigma_{f}^{\star})$
This paper Average iterate, $\mathrm{diam}(\mathcal{X})\leq D$ AdaGrad-Norm, AdaFTRL Yes $\mathcal{O}(\tfrac{1}{T}+\tfrac{1}{\sqrt{T}}\sigma_{g}^{\star})$
@@
As a byproduct, we are also able to slightly refine the existing convergence analysis of AdaGrad-Norm for unconstrained smooth convex optimization [20].
@@
4.2 Constraint-aware $G^{\star}$ regret via smooth surrogate loss
Although this paper primarily focuses on constrained OCO, the definition of $L^{\star}_{T}$ measures suboptimality over $\mathbb{R}^{n}$ and does not take into account the constraint set $\mathcal{X}$. A more desirable measure is
$L^{\mathcal{X}}_{T}(x)\coloneqq\sum_{t=1}^{T}\ell_{t}(x)-\min_{u\in\mathcal{X}}\ell_{t}(u)$
rather than $\sum_{t=1}^{T}\ell_{t}(x)-\inf\ell_{t}$ (or $\sum_{t=1}^{T}\ell_{t}(x)$ for non-negative losses). A similar observation applies to $G^{\star}$: the gradient norm $\|\nabla\ell_{t}(x)\|$ does not incorporate the constraint set $\mathcal{X}$, whereas a more appropriate measure should involve $\mathrm{dist}(-\nabla\ell_{t}(x),\mathcal{N}_{\mathcal{X}}(x))$, where $\mathcal{N}_{\mathcal{X}}(x)\coloneqq\{g:\langle g,x-y\rangle\geq 0,\text{ for all }y\in\mathcal{X}\}$ denotes the normal cone of $\mathcal{X}$ at $x$. To illustrate the benefit of $L^{\mathcal{X}}_{T}$, suppose there exists $x^{\star}\in\operatorname{bd}\mathcal{X}$ that is the constrained optimum for most losses. In this case, $L^{\mathcal{X}}_{T}(x^{\star})=\mathcal{O}(1)$, whereas $L^{\star}_{T}$ can still be $\Omega(T)$ if the unconstrained optima are excluded by $\mathcal{X}$. Incorporating constraints into small loss (and gradient norm) measures is therefore advantageous. While there have been many successful extensions of OCO theory that better exploit constraint geometry (for example, online mirror descent [24] and AdaGrad [8]), these approaches are not fully compatible with the $L^{\star}$ or $G^{\star}$ regret. In the remainder of this section, we illustrate a strategy that enables dependence on $L^{\mathcal{X}}_{T}$ and a corresponding stationarity measure by replacing the losses with suitable smooth surrogates.
Smooth approximation.
Given a lower-semicontinuous convex function $\ell:\mathcal{X}\rightarrow\mathbb{R}$, its associated Moreau envelope with parameter $\gamma>0$, denoted by $\ell^{1/\gamma}(x)$ [3], is given by
@@ 4.2 Constraint-aware $G^{\star}$ regret via smooth surrogate loss
Regardless of the domain or smoothness of $\ell$, the funcition $\ell^{1/\gamma}$ is always $\gamma$-smooth and convex, with gradient given by $\nabla\ell^{1/\gamma}(x)=\gamma(x-\operatorname{prox}_{\ell/\gamma}(x))$, where $\operatorname{prox}_{\ell/\gamma}(x)=\operatornamewithlimits{arg\,min}_{x}\{\ell(y)+\tfrac{\gamma}{2}\|y-x\|^{2}\}$ is the proximal point. In particular, $\ell^{1/\gamma}$ and $\ell$ share the same minimizers and stationarity points, and $\nabla\ell^{1/\gamma}(x)$ is an approximation to the quantity $\mathrm{dist}(-\nabla\ell_{t}(x),\mathcal{N}_{\mathcal{X}}(x))$ [6]. Hence, we similarly define $G^{\mathcal{X}}_{T}(x)\coloneqq\sum_{t=1}^{T}\|\nabla\ell^{1/\gamma}(x)\|^{2}$. In view of the properties of the smooth approximation, instead of solving the original OCO problem, we consider the OCO problem defined on the smooth surrogates:
@@
Since each loss $\ell^{1/\gamma}_{t}$ is convex and $\gamma$-smooth, the results from the previous sections apply.
@@
We have $G^{\mathcal{X}}_{T}(x)\leq 2\gamma L^{\mathcal{X}}_{T}(x)$, and all the results from previous sections hold for $\rho^{1/\gamma}_{T}$. In particular, there exist online algorithms satisfying $\rho_{T}^{1/\gamma}(x)=\mathcal{O}(\sqrt{G^{\mathcal{X}}_{T}(x)})=\mathcal{O}(\sqrt{\gamma L^{\mathcal{X}}_{T}(x)})$ for any $x\in\mathcal{X}$.
We make two remarks on running online learning algorithms on the surrogate loss sequence $\{\ell^{1/\gamma}_{t}\}_{t=1}^{T}$. First and importantly, the regret guarantee in Proposition˜4.1 is with respect to $\rho_{T}^{1/\gamma}$ instead of $\rho_{T}$, and the gap between $\ell^{1/\gamma}(x^{t})$ and $\ell(x^{t})$ can be up to $\mathcal{O}(\gamma^{-1})$ [3]. Therefore, $\gamma=\Omega(\sqrt{T})$ is required to ensure $\rho_{T}(x)\leq\rho_{T}^{1/\gamma}(x)+\mathcal{O}(\sqrt{T})$, and there is no regret improvement in the worst-case. Nevertheless, whenever it is feasible to directly redefine the loss, using the smoothed loss enables adaptivity to the constrained small loss bound $L^{\mathcal{X}}_{T}(x)$.
Second, running online gradient descent on $\ell^{1/\gamma}_{t}$ requires finding $\nabla\ell_{t}^{1/\gamma}$, which requires a proximal point iteration and can often be efficiently evaluated [29]. Note that running an online algorithm on smooth surrogates $\{\ell^{1/\gamma}_{t}\}_{t=1}^{T}$ differs from the existing implicit online learning (online proximal point) approach [19], since proximal point is applied not as an online algorithm itself, but instead as a subroutine to find the envelope gradient $\nabla\ell_{t}^{1/\gamma}$. This smoothing strategy can be applied jointly with any existing adaptive online algorithm, such as AdaGrad.

### v3

@@ 3.1 Dynamic $G^{\star}$ regret
Section˜2 focuses on the static regret with respect to a single competitor $x\in\mathcal{X}$. In this subsection, we consider dynamic regret, which allows for a sequence of competitors at the cost of a path length term in the final regret bound. Given a comparator sequence $\{\hat{x}_{t}\}_{t=1}^{T}$ with path length $\mathcal{P}_{T}(\{\hat{x}_{t}\}_{t=1}^{T})\coloneqq\sum_{t=2}^{T}\|\hat{x}_{t}-\hat{x}_{t-1}\|$, the dynamic regret with respect to this sequence is defined as
@@ 3.2 $G^{\star}$ regret under bandit feedback
This smoothed function $\tilde{\ell}_{t}$ naturally serves as a surrogate function when performing BGD. Key properties of the smoothed function used in our analysis are summarized in the appendix. The following theorems establish regret bounds for BGD under both constant and adaptive learning rates.
@@
Under A1, A2 and suppose $n\geq 8$, bandit gradient descent (Algorithm˜5 in the appendix) $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta g_{t}]$ with constant learning rate $\eta\in(0,\tfrac{1}{4nL})$ satisfies
$\mathbb{E}[\rho_{T}(x)]\leq\textstyle\tfrac{D^{2}}{2\eta}+\tfrac{4n\eta}{1-4n\eta L}G_{T}(x)+\tfrac{4n\eta}{1-4n\eta L}TL^{2}\mu^{2}+\tfrac{\eta T}{2}n^{2}L^{2}\mu^{2}+\tfrac{TL\mu^{2}}{2}$
for any $x\in\mathcal{X}$. In particular, choosing $x=x^{\star}$, $\eta=\min\big\{\tfrac{D}{4\sqrt{nG^{\star}_{T}}},\tfrac{1}{8nL}\big\}$ and $\mu=\tfrac{D}{\sqrt{2nT}L}$ gives $\mathbb{E}[\rho_{T}(x^{\star})]\leq\max\{8nLD^{2},4D\sqrt{nG^{\star}_{T}}\}+\tfrac{LD^{2}}{2}$.
@@ 4.1 Stochastic optimization under interpolation
$(\sigma_{f}^{\star})^{2}\coloneqq\inf_{x}\mathbb{E}[f(x,\xi)-\inf_{u}f(u,\xi)]\quad\text{and}\quad(\sigma_{g}^{\star})^{2}\coloneqq{\mathbb{E}[\|\nabla f(x^{\star},\xi)\|^{2}]},$
where $x^{\star}\in\operatornamewithlimits{arg\,min}_{x}f(x)$ is assumed to exist. We make the following simplifying assumption.
@@ 1. A3:
Under A3, we have $(\sigma_{f}^{\star})^{2}=\min_{x\in\mathcal{X}}\mathbb{E}[f(x,\xi)-\inf_{u}f(u,\xi)]$, and it is easy to verify that $\sigma_{f}^{\star}$ and $\sigma_{g}^{\star}$ mirror the $L^{\star}$ and $G^{\star}$ regret, respectively. Let $\{\xi^{t}\}_{t=1}^{T}$ be i.i.d. samples from distribution $\Xi$. Taking $\ell_{t}(x)=f(x,\xi^{t})$ gives
$\mathbb{E}[L_{T}^{\star}]={}$ $\textstyle\mathbb{E}[\sum_{t=1}^{T}\ell_{t}(x^{\star})-\inf\ell_{t}]=\mathbb{E}[\sum_{t=1}^{T}f(x^{\star},\xi^{t})-\inf f(u,\xi^{t})]=T(\sigma_{f}^{\star})^{2},$
$\mathbb{E}[G_{T}^{\star}]={}$ $\textstyle\mathbb{E}[\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{\star})\|^{2}]=\mathbb{E}[\sum_{t=1}^{T}\|\nabla f(x^{\star},\xi^{t})\|^{2}]=T(\sigma_{g}^{\star})^{2}.$
When $\sigma_{f}^{\star}$ or $\sigma_{g}^{\star}$ is close to 0, the regime is known as (near-)interpolation [22]. Interpolation is a central setting in optimization for machine learning and has recently attracted significant attention [29, 21, 33, 2]. Since online algorithms with sublinear regret can be converted into an optimization algorithm via online-to-batch conversion [25], our $G^{\star}$ regret bounds from Section˜2.2 imply convergence guarantees for stochastic gradient methods and stochastic dual averaging (FTRL for linearized losses).
@@
  • if OGD in Theorem˜2.1 is applied to $\{\ell_{t}\}_{t=1}^{T}$, then for any $\eta\in(0,\tfrac{1}{L})$, we have
@@
  • if AdaGrad-Norm in Theorem˜2.2 is applied to $\{\ell_{t}\}_{t=1}^{T}$, then $\mathbb{E}[f(\bar{x}^{T})-f(x^{\star})]\leq\tfrac{LD^{2}}{T}+\tfrac{\sqrt{2}D}{\sqrt{T}}\sigma_{g}^{\star}$;
@@
Corollary˜4.1 improves the existing results [29] by replacing $\sigma_{f}^{\star}$ by $\sigma_{g}^{\star}$. Compared with the recent last-iterate convergence results of stochastic gradient descent [2, 12], our result enables adaptivity to smoothness and noise for the average iterate (Table˜1). Furthermore, using the anytime online-to-batch conversion from [5], similar last-iterate convergence follows immediately.
Corollary 4.2.
Under the same assumptions as Corollary˜4.1, algorithms in Corollary˜4.1 with anytime online-to-batch conversion [5] give the same guarantees for the last iterate $x^{T}$.
Table 1: Implications of the $G^{\star}$ regret for stochastic smooth convex optimization. Prefix AT denotes algorithm variants with anytime online-to-batch conversion [5]. The settings with bounded domain also assume that A3.
@@
[29] Avg. iter, $\mathcal{X}=\mathbb{R}^{n}$ SGD No $\mathcal{O}(\frac{1}{T}+\frac{\sqrt{L}}{\sqrt{T}}\sigma_{f}^{\star})$
[2, 12] Last iter, $\mathcal{X}=\mathbb{R}^{n}$ SGD No $\mathcal{O}\big(\frac{\log T}{T}+\sqrt{\frac{\log T}{T}}\,\sigma_{g}^{\star}\big)$
[32] Avg. iter, $\mathrm{diam}(\mathcal{X})\leq D$ AdaGrad Yes $\mathcal{O}(\frac{1}{T}+\frac{\sqrt{L}}{\sqrt{T}}\sigma_{f}^{\star})$
This paper Avg. iter, $\mathrm{diam}(\mathcal{X})\leq D$ AdaGrad-Norm Yes $\mathcal{O}(\frac{1}{T}+\frac{1}{\sqrt{T}}\,\sigma_{g}^{\star})$
AdaFTRL
Last iter, $\mathrm{diam}(\mathcal{X})\leq D$ AT-AdaGrad-Norm Yes $\mathcal{O}(\frac{1}{T}+\frac{1}{\sqrt{T}}\,\sigma_{g}^{\star})$
AT-AdaFTRL
@@
As a byproduct, we also slightly refine the existing analysis of AdaGrad-Norm for unconstrained smooth convex optimization [20].
@@
4.2 (Towards) constraint-aware $L^{\star}$ (and $G^{\star}$) regret via smooth surrogate loss
Although this paper focuses on constrained OCO, the definition of $L^{\star}_{T}$ measures suboptimality over $\mathbb{R}^{n}$ and does not take into account the constraint set $\mathcal{X}$. A more desirable measure is
$L^{\mathcal{X}}_{T}(x)\coloneqq\sum_{t=1}^{T}\ell_{t}(x)-\min_{u\in\mathcal{X}}\ell_{t}(u),$
which measures suboptimality relative to the constrained optimum rather than the unconstrained optimum $\sum_{t=1}^{T}\ell_{t}(x)-\inf\ell_{t}$. Similarly for $G^{\star}$, the gradient norm $\|\nabla\ell_{t}(x)\|$ is oblivious of the constraint set $\mathcal{X}$. Instead, a more appropriate measure should involve $\mathrm{dist}(-\nabla\ell_{t}(x),\mathcal{N}_{\mathcal{X}}(x))$, where $\mathcal{N}_{\mathcal{X}}(x)\coloneqq\{g:\langle g,x-y\rangle\geq 0,\text{ for all }y\in\mathcal{X}\}$ denotes the normal cone of $\mathcal{X}$ at $x$. The constrained cumulative suboptimality $L^{\mathcal{X}}_{T}$ can provide sharper problem-dependent constants. For example, suppose the unconstrained optimum is not in $\mathcal{X}$ but the constrained optimum $x^{\star}\in\operatorname{bd}\mathcal{X}$ is also the constrained optimum for most losses individually: $\ell_{t}(x^{\star})=\min_{x\in\mathcal{X}}\ell_{t}(x)$. In this case, $L^{\mathcal{X}}_{T}(x^{\star})=\mathcal{O}(1)$, whereas $L^{\star}_{T}$ can still be $\Omega(T)$. The rest of this section demonstrates the (im)possibility of generalizing the existing small loss (gradient) bound to adapt to the constraint set $\mathcal{X}$.
Hardness.
As discussed, the naive goal is to design an online algorithm that directly adapts to $L^{\mathcal{X}}_{T}$ (or $G^{\mathcal{X}}_{T}$) by replacing the traditional $\mathcal{O}(LD^{2}+D\sqrt{LL^{\star}_{T}})$ by $\mathcal{O}(LD^{2}+D\sqrt{LL^{\mathcal{X}}_{T}(x^{\star})})$. However, Proposition˜4.1 shows that this goal is generally not achievable.
Proposition 4.1.
For any $\tau\in[0,\tfrac{1}{6})$, there exists a loss sequence $\{\ell_{t}\}_{t=1}^{T}$ satisfying A1, A2 and that $\inf\ell_{t}=0$ such that for any online algorithm and $x\in\mathcal{X}$, $\frac{\rho_{T}(x)}{LD^{2}+D\sqrt{LL^{\mathcal{X}}_{T}(x)}}=\Omega(T^{\tau})$.
Despite the hardness result, whenever it is possible to redefine the losses, we show a strategy that enables dependence on $L^{\mathcal{X}}_{T}$ and $G^{\mathcal{X}}_{T}$ by replacing the losses by their smooth surrogates.
Easiness with smoothing.
Given a lower-semicontinuous convex function $\ell:\mathcal{X}\rightarrow\mathbb{R}$, its associated Moreau envelope with parameter $\gamma>0$, denoted by $\ell^{1/\gamma}(x)$, is given by
@@
Regardless of the domain or smoothness of $\ell$, the function $\ell^{1/\gamma}$ is always $\gamma$-smooth and convex, with gradient given by $\nabla\ell^{1/\gamma}(x)=\gamma(x-\operatorname{prox}_{\ell/\gamma}(x))$, where $\operatorname{prox}_{\ell/\gamma}(x)=\operatornamewithlimits{arg\,min}_{x}\{\ell(y)+\tfrac{\gamma}{2}\|y-x\|^{2}\}$ is the proximal point. In particular, $\ell^{1/\gamma}$ and $\ell$ share the same minimizers and stationarity points, and $\nabla\ell^{1/\gamma}(x)$ approximates the quantity $\mathrm{dist}(-\nabla\ell(x),\mathcal{N}_{\mathcal{X}}(x))$ [7]. Hence, we similarly define $G^{\mathcal{X}}_{T}(x)\coloneqq\sum_{t=1}^{T}\|\nabla\ell_{t}^{1/\gamma}(x)\|^{2}$. In view of the properties of the smooth approximation, instead of solving the original OCO problem, suppose we consider the OCO problem defined on the smooth surrogates:
@@
Since each loss $\ell^{1/\gamma}_{t}$ is convex and $\gamma$-smooth, the results from the previous sections directly apply.
@@
We have $G^{\mathcal{X}}_{T}(x)\leq 2\gamma L^{\mathcal{X}}_{T}(x)$, and the results from previous sections hold for $\rho^{1/\gamma}_{T}$. In particular, there exist online algorithms satisfying $\rho_{T}^{1/\gamma}(x)=\mathcal{O}(\gamma D^{2}+\sqrt{G^{\mathcal{X}}_{T}(x)})=\mathcal{O}(\gamma D^{2}+\sqrt{\gamma L^{\mathcal{X}}_{T}(x)})$ for any $x\in\mathcal{X}$.
We make two remarks on running online learning algorithms on the surrogate loss sequence $\{\ell^{1/\gamma}_{t}\}_{t=1}^{T}$.
First and importantly, the regret guarantee in Proposition˜4.2 is with respect to $\rho_{T}^{1/\gamma}$ instead of $\rho_{T}$, and the gap between $\ell^{1/\gamma}(x^{t})$ and $\ell(x^{t})$ can be up to $\mathcal{O}(\gamma^{-1})$ [3, 7]. In particular, using Proposition 4.2 of [7], we have
$\ell^{1/\gamma}(x)+\tfrac{1}{2\gamma}\|\nabla\ell^{1/\gamma}(x)\|^{2}\leq\ell(x)\leq\ell^{1/\gamma}(x)+\tfrac{1}{\gamma}\|\nabla\ell(x)\|^{2},$ (16)
and combining (16) with Proposition˜4.2 yields a bound on the regret:
$\rho_{T}(x)\leq\rho^{1/\gamma}_{T}(x)+\frac{1}{\gamma}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}\leq\mathcal{O}\big(\gamma D^{2}+\frac{1}{\gamma}\sum_{t=1}^{T}\|\nabla\ell_{t}(x^{t})\|^{2}+\sqrt{G_{T}^{\mathcal{X}}(x)}\big).$ (17)
Assuming $\|\nabla\ell(x^{t})\|\leq M$, the first two terms $\gamma D^{2}+\frac{1}{\gamma}TM^{2}$ require $\gamma=\Omega(\sqrt{T})$ to guarantee $\rho_{T}(x)\leq\mathcal{O}(\sqrt{T})$, and there is no regret improvement in the worst-case. In the appendix, we show the hardness of obtaining an $\mathcal{O}(LD^{2}+D\sqrt{G^{\mathcal{X}}_{T}(x)})$ on $\rho_{T}(x)$ using an argument similar to Proposition˜4.1. While this argument contains an improved bound for the smoothed loss, it does not always yield an improved bound for the original loss.
Remark 7.
The $G^{\star}$ regret can be strictly tighter than the $L^{\star}$ regret. To see this, notice that for $\gamma=\Theta(\sqrt{T})$, the $\mathcal{O}(\gamma D^{2}+\sqrt{\gamma L^{\mathcal{X}}_{T}(x)})$ upper bound yields a $\sqrt{\gamma L^{\mathcal{X}}_{T}(x)}$ term that can be as large as $\Theta(T^{3/4})$ in the worst case.
Second, running OGD on $\ell^{1/\gamma}_{t}$ requires computing $\nabla\ell_{t}^{1/\gamma}$. This quantity can be evaluated using the proximal point iteration, which is often efficient [28]. Note that running an online algorithm on smooth surrogates $\{\ell^{1/\gamma}_{t}\}_{t=1}^{T}$ differs from the existing implicit online learning (proximal point) approach [19], since proximal point is applied not as an online algorithm itself, but instead as a subroutine to find the envelope gradient $\nabla\ell_{t}^{1/\gamma}$. This smoothing strategy can be applied jointly with any existing adaptive online algorithm, such as AdaGrad.